Consider equally risky, all-equity financed firms G and D. Both firms are currently trading at $50 per share. Firm G pays no dividends, but firm D pays all its earnings as dividends. In a year from today, the stock of firm G is expected to be at $65 per share. Firm D is expected to pay a $15 dividend per share at the year-end and its stock price at the year-end is expected to be $50 per share. Capital gains are taxed at half the 40% regular income tax rate, but the dividend income is taxed at 30%.

Required At what share price should the stock of firm D be trading to-day in a competitive stock market? Show calculations to support your answer.
Answer #1

capital gains are taxed at 20% (40%/2)

capital gains on firm G share = 65 - 50 = 15 (price after one year - current price)

tax = 15*20% = 3

after tax return = 15 - 3 = 12

return in % = (after tax return / current price) * 100

= (12 / 50) * 100

= 24%

to stay in competitive stock D should provide after tax return of 24%

tax on dividend = 30%

Dividend = 15

after tax dividend = 15*(1-0.3) = 10.50

return = (after tax dividend / Price)*100

24% = (10.50 / Price) * 100

so Price should be = $43.75

(please give thumbs up if it's helpful.thank you)